OpenStudy (anonymous):

Evaluate indefinite integral of ∫[v+(1/(3v-1)^3)]dv.

OpenStudy (raden):

∫[v+(1/(3v-1)^3)]dv = ∫v dv + ∫[1/(3v-1)^3]dv ∫v dv = ... ( it's easy :) ∫[1/(3v-1)^3]dv use integral by u-sub u = 3v - 1 du = 3 dv or dv = du/3 continue it ...